ELECTRONIC SPECTRUM OF THE AlC_{2} RADICAL

\underline{EVAN B. JOCHNOWITZ}, EGOR CHASOVSKIKH, EUNSOOK KIM, and JOHN P. MAIER, Department of Chemistry, University of Basel, Klingelbergstrasse 80, CH-4056 Basel, Switzerland; ISABELLE NAVIZET, Laboratoire de Chimie Théorique, Université de Marne-la-Vallée, F-77454 Champs sur Marne, France.

An electronic transition of the AlC_{2} radical (C_{2v} structure) has been observed using laser induced fluorescence spectroscopy. The molecule was prepared in a supersonic expansion using laser ablation of an aluminum rod in the presence of acetylene gas. A spectrum was recorded and assigned to the C̃ ^{2}B_{2} - X̃ ^{2}A_{1} system based on a rotational analysis and agreement with calculated molecular parameters and excitation energies. Ab initio results are in accord with previous theoretical studies which conclude that AlC_{2} possesses a triangular C_{2v} ^{2}A_{1} geometry, with the linear C_{∞v} ^{2}Σ^{+} AlCC isomer 0.70 eV higher in energy. A spectral fit yields molecular constants both in the ground and electronically excited states.